OpenStudy (anonymous):

Solve Each Equation by first clearing the fractions or decimals 0.2(c-4)+1=0.15(c+2) I Need to show all work. I Am Horrible At Math Can Someone Please Take Me Through The Steps to get Positive 2 For The Answer

OpenStudy (anonymous):

first multiply both sides by 100, which is the same as moving the decimal two places to the left. you get \[20(c-4)+100=15(c+2)\]

OpenStudy (anonymous):

then remove the parentheses using the distributive law you get \[20c-80+100=15c+30\] and since \(100-80=20\) this becomes \[20c+20=15c+30\]

OpenStudy (anonymous):

subtract \(15c\) from both sides, gives \[5c+20=30\]

OpenStudy (anonymous):

subtract 20 from both sides gives \[5c=10\] then divide both sides by \(5\) and get \[c=10\div 5=0\]